Heights of the students in a class are given in the distribution below:



5' - 5'2"      10



5'2" - 5'4" 40



    5'4" - 5'6"   25 .



Find:



1. Central Tendency



2. Dispersion



3. Skewness and kurtosis





1
Expert's answer
2022-03-17T03:45:51-0400

Solution

1. Central tendency is the measurement of mean, median and the mode.




(a) Mean.

"\bar X=\dfrac{\sum fX}{\sum f}"


"=\dfrac{400.5}{75}=5.34"


(b) median

"\dfrac{N}{2}=\dfrac{75}{2}=37.5"


Hence the median lies in the class "5.2-5.4"


"Median=L+(\dfrac{\frac{N}{2}-pcf}{f})i"


"=5.2+(\dfrac {37.5-11}{40})0

.2"


"=5.3325"


Hence the median height is "5'3"


(c) Mode

The class with the highest frequency "(40)" is "5.2-5.4" and hence is the modal class.


"Mode=L(\dfrac{\Delta_1}{\Delta_1+\Delta_2})i"


"M_0=5.2(\dfrac{(40-10)}{(40-10)+(40-25)})0.2"


"=5.333"


Modal height is "5'3""


2. Dispersion is the measurement of variance and standard deviation.





(a) Variance

"\sigma^2=\dfrac {\sum f(X-mean)^2}{n-1}"


"\sigma^2=\dfrac{1.28}{74}=0.017"


(b) Standard deviation

"\sigma=\sqrt {variance}=\sqrt{0.017}"


"\sigma=0.132"


3. (a) Skewness


"=\dfrac{\sum (X-mean)^3}{n.\sigma^3}"


"=\dfrac{-0.009792}{75\times0.132^3}=-0.057"


"=-0.057"


(b) Kurtosis

"=\dfrac{\sum(X-mean)^4}{n.\sigma^4}"


"=\dfrac{0.00397568}{75\times0.132^4}=0.175"


"=0.175"